# Spacetime foam correlation renders cosmological constant (dark energy)

### v1

@@ 4.1 Infrared fixed point and scaling invariant domain
[…]
@@ 4.4 Ultraviolet fixed point and scaling invariant domain
[…]
@@
[…]
@@ 5.1 Effective action for field operators at large distances
Here we resume the subscript “$i$” for types of foamon field. In terms of the effective Lagrangian ${\mathcal{L}}_{\rm eff}$ (4.2) and (4.27) in the scaling invariant domain, the Euclidean generating function (3.7) becomes
@@ 5.3 Nonlocal field action originated from spacetime foam correlations
This is a non-local operator, since the foamon field $\phi_{i}$ propagation $(\Delta_{ij})^{-1}=\delta_{ij}(\partial^{\prime 2}+m^{2}_{i})^{-1}$ couples two operators $O_{i}$ at different spacetime points. The two-point function (2.11) is exponentially suppressed for $|x^{\prime}_{2}-x^{\prime}_{1}|>(m_{i})^{-1}$, indicating that $O_{i}(x^{\prime}_{1})$ and $O^{\dagger}_{i}(x^{\prime}_{2})$ are not correlated if they are separated beyond $(m_{i})^{-1}$. It measures the size of the operator $O_{i}$ locality.
@@
[…]
@@
By the Wick rotation on the effective Euclidean action (5.15), we obtain the effective action $S^{\rm eff}_{M}(\tilde{O}_{i})$ in the Minkowski space.
@@ 6.2 Equation of state and interaction with matter
As the Universe volume $V\propto H^{-3}$ adiabatically expands $\delta V>0$, conserving the total number of degrees of freedom $\delta N=0$, the positive particle’s repulsive energy $E$ (mean kinetic energy) decreases $\delta E<0$, i.e., energy density $\rho$ and pressure $p$ decreases. It obeys the first law of thermodynamics in the comoving frame,
[…]
@@
The foamon field energy density (6.2) can be formally written as $\rho_{{}_{\Lambda}}=M_{\rm pl}n_{w}$ in terms of the effective foam number density $n_{w}=\ell^{-1}_{\rm pl}\xi^{-2}/(8\pi)^{2}$, as if the relevant modes of the spacetime foamon field were holographically distributed as a thin layer of width $4\pi\xi^{2}\ell_{\rm pl}$ near to the Universe horizon $\xi\sim H^{-1}$. The holographical configuration minimises the total energy and number of spacetime foamon modes, whose wavelength is the order of the Universe horizon. The possible interpretation is that in the macroscopic average, the local foamon field fluctuations (oscillations) in the homogeneous space cancel each other up to the horizon. Here, the local foamon field oscillation exhibits similar dynamics to the “microcyclic universes” at small scales, shown by local scale factor oscillation in Figure 1 of Refs. [32, 33].
@@ 6.3 Scale factor and renormalisation group flow
In the Friedmann–Lemaître–Robertson–Walker (FLRW) cosmology, the scale factor $a(t)$ stretches the physical spacetime $x_{\rm phys}=a(t)x_{\rm com}$ and $k_{\rm phys}=k_{\rm com}/a(t)$, where $x_{\rm com}$ and $k_{\rm com}$ are in the comoving coordinate. On the physical spacetime, the renormalised foam fields, masses, and couplings of the effective Lagrangian (4.1) or (4.25) should correspond to observables at long distances. It implies that as the physical spacetime length scale stretches in the universe expansion, the small-scale foamon field modes are integrated, and renormalised fields, masses, couplings and effective Lagrangian (4.2) should follow the Wilson renormalisation group equations in the scaling invariant domain of physical infrared or ultraviolet fixed point. From this point of view, we identify the correspondences between the scaling parameter $b$ (3.17) and the scalar factor $a(t)$,
@@
The infrared fixed point $b\sim 0$ corresponds to $a(t)\sim 0$ for the early Universe. The ultraviolet fixed point $b\sim 1$ corresponds to $a(t_{0})\sim 1$ for the present $t_{0}$ epoch of the Universe.
@@ 7 Summary and remarks
We end this article with a few speculations in this scenario. The spatial distribution of the foamon field energy, the first term in (5.15), should not be homogeneous. In a galaxy, the foamon field energy relates to the Ricci scalar scale that depends on galactic matter. It is a positive energy bound inside a negative gravitational potential. From the viewpoint of galactic stability, it must affect the galaxy’s rotation curve, and its concave gravitational lensing effects should not be negligible compared with matter’s convex ones. In the sector of the foamon field $\Phi_{f}$ associated with the fermion operator $O_{f}=(\bar{\psi}_{L}\psi_{R})$, from the effective action (5.15), we obtain the effective four-fermion operator of Einstein-Cardan type

### v2

@@ 4.1 Infrared fixed point and scaling invariant domain
The $(\cdot\cdot\cdot)$ indicates the counterterms and irrelevant high-dimensional operators suppressed by ${\mathcal{O}}(b^{l})$ and $l\geq 1$. In the scaling invariant domain, the field $\phi$ wave-function, mass and coupling corrections $\delta Z$, $\delta m^{2}$ and $\delta\lambda$, as well as irrelevant operators $(\cdot\cdot\cdot)$ can be ignored. It gives the Wilson scaling invariance $\Gamma(\phi)\approx\Gamma(\phi_{\rm ir})$ with $\phi\approx b^{\frac{2-d}{2}}\phi_{\rm ir}$ and $m^{2}\approx m_{\rm ir}^{2}b^{-2}$ and $\lambda\approx\lambda_{\rm ir}b^{d-4}$, because the correlation length $\xi=1/m_{\rm ir}$ is much larger than the scale variation (3.17).
@@ 4.4 Ultraviolet fixed point and scaling invariant domain
The $(\cdot\cdot\cdot)$ represents the counterterms and high-dimensional $(d>4)$ operators. The scalar field $\phi_{\rm uv}$ becomes massive, and massless Goldstone bosons $\pi_{\rm uv}$ become nonlinear Sigma model or longitudinal modes of massive gauge bosons coupling to fermions, which will be discussed in a separate article.
@@
when $b$ increases. All irrelevant operators $(\cdot\cdot\cdot)$ are suppressed by ${\mathcal{O}}(b^{l})$ and $l>1$. The effective theory is the Wilson scaling invariant $\Gamma(\phi,\psi)\approx\Gamma(\phi_{\rm uv},\psi_{\rm uv})$ with the correlation length $\xi=1/m_{\rm uv}$.
@@ 5.1 Effective action for field operators at large distances
Here we resume the subscript “ii” for types of foamon field. In terms of the effective Lagrangian ${\mathcal{L}}_{\rm eff}$ (4.2) and (4.27) in the scaling invariant domain, the Euclidean generating function (3.7) becomes
@@ 5.3 Nonlocal field action originated from spacetime foam correlations
This is a non-local operator, since the foamon field $\phi_{i}$ propagation $(\Delta_{ij})^{-1}=\delta_{ij}(\partial^{\prime 2}+m^{2}_{i})^{-1}$ couples two operators $O_{i}$ at different spacetime points. This effective nonlocal action is attributed to the correlation functions (2.10) and (2.12) of the foamon fields in different spacetime points, in contrast with the connection function $C_{ij}(x_{1},x_{2})$ (1.1) of wormholes connecting two spacetime points. The two-point function (2.11) is exponentially suppressed for $|x^{\prime}_{2}-x^{\prime}_{1}|>(m_{i})^{-1}$, indicating that $O_{i}(x^{\prime}_{1})$ and $O^{\dagger}_{i}(x^{\prime}_{2})$ are not correlated if they are separated beyond $(m_{i})^{-1}$. It measures the size of the operator $O_{i}$ locality.
@@
The high-dimensional derivative operators $(\cdot\cdot\cdot)$ represent nonlocal contributions, which are suppressed for $b<1$ in the scaling domains (4.2) and (4.27). We will henceforth neglect the nonlocal contributions $(\cdot\cdot\cdot)$. Thus, in Eq. (5.9) we approximately obtain the local operator at the leading order,
@@
By the Wick rotation on the effective Euclidean action (5.15) at long distances, we obtain the corresponding effective action $S^{\rm eff}_{M}(\tilde{O}_{i})$ in the Minkowski space. We emphasise that the Wick rotation acts on the effective Euclidean action (2.3) and path integral (3.3) of the foamon field $\Phi_{i}$ coupling to the field operator $O_{i}$ at large distances, rather than the Euclidean gravity action $I_{E}$ (1.3) and path integral $\int{\mathcal{D}}ge^{-I_{E}}$.
@@ 6.2 Equation of state and interaction with matter
As the Universe volume $V\propto H^{-3}$ adiabatically expands $\delta V>0$, conserving the total number of degrees of freedom $\delta N=0$, the positive particle’s repulsive energy $E$ (mean kinetic energy) decreases $\delta E<0$, i.e., energy density $\rho$ and pressure $p$ decrease. It obeys the first law of thermodynamics in the comoving frame,
$p\delta V+\delta E=0,\quad p=-(\delta E/\delta V)>0,\quad E=\int dV\rho.$ (6.3)
@@
The foamon field energy density (6.2) can be formally written as $\rho_{{}_{\Lambda}}=M_{\rm pl}n_{w}$ in terms of the effective foam number density $n_{w}=\ell^{-1}_{\rm pl}\xi^{-2}/(8\pi)^{2}$, as if the relevant modes of the spacetime foamon field were holographically distributed as a thin layer of width $4\pi\xi^{2}\ell_{\rm pl}$ near to the Universe horizon $\xi\sim H^{-1}$. The holographic configuration minimises the total energy and number of spacetime foamon modes, whose wavelength is of the order of the Universe horizon. The possible interpretation is that in the macroscopic average, the local foamon field fluctuations (oscillations) in the homogeneous space cancel each other up to the horizon. Here, the local foamon field oscillation exhibits similar dynamics to the “microcyclic universes” at small scales, shown by local scale factor oscillation in Figure 1 of Refs. [44, 45].
@@ 6.3 Scale factor and renormalisation group flow
In the Friedmann–Lemaître–Robertson–Walker (FLRW) cosmology, the scale factor $a(t)$ stretches the physical spacetime $x_{\rm phys}=a(t)x_{\rm com}$ and $k_{\rm phys}=k_{\rm com}/a(t)$, where $x_{\rm com}$ and $k_{\rm com}$ are in the comoving coordinate. On the physical spacetime, the renormalised foam fields, masses, and couplings of the effective Lagrangian (4.1) or (4.25) should correspond to observables at long distances. It implies that as the physical spacetime length scale stretches in the universe expansion, the small-scale foamon field modes are integrated, and renormalised fields, masses, couplings and effective Lagrangian (4.2) should follow the Wilson renormalisation group equations in the scaling invariant domain of physical infrared or ultraviolet fixed point. From this point of view, we identify the correspondences between the scaling parameter $b$ (3.17) and the scale factor $a(t)$,
@@
The infrared fixed point $b\sim 0$ corresponds to $a(t)\sim 0$ for the early Universe. The ultraviolet fixed point $b\sim 1$ corresponds to $a(t_{0})\sim 1$ for the present $t_{0}$ epoch of the Universe. The scaling parameter $b$ plays the same role of the scale factor $a(t)$ in a stretching FLRW Universe.
@@ 7 Summary and remarks
We end this article with a few speculations on this scenario for dark energy. In local universes, the matter distributions are not homogeneous, and the dark energy distributions should not be homogeneous. The reason is that the foamon field dark energy, the first term in (5.15), depends on the Ricci scalar $R$ of self-gravitational systems of matter and dark energy, which obey the Einstein equation. The induced matter and dark energy interactions result in matter and dark energy inhomogeneous distributions in an equilibrium configuration. The foamon field dark energy in a galaxy relates to the Ricci scalar, which depends on galactic matter. It is a positive repelling energy bound inside a negative attractive potential due to self-gravitating matter. Thus, the interactions between matter and dark energy establish the gravitational potentials for stable galactic systems. Consequently, these interactions and resultant gravitational potential must lead to the flattened galaxy’s rotation curve and the smoothed gravitational lensing effects, since the contributions from matter and dark energy are opposite. These dynamics and effects deserve studies. We note the studies [56, 57] using phenomenological models for dark energy.
In the sector of the foamon field $\Phi_{f}$ associated with the fermion operator $O_{f}=(\bar{\psi}_{L}\psi_{R})$, from the effective action (5.15), we obtain the effective four-fermion operator of Einstein-Cardan type